### v1

@@ 3.4.2 Conditional Density Estimation
A key step in this process involves estimating the conditional density function of the intervention given the confounders. Accurate conditional density estimation is critical because it helps quantify the regions where the oracle intervention is supported by the data. Various methods can be employed for this task, including parametric models, non-parametric kernel conditional density estimation 12, and more flexible approaches like the highly adaptive lasso density estimation (Haldensify) 13, 14, 15. Inspired by Haldensify, in this paper we use a specific hazard binning conditional density estimation procedure, which we describe in Appendix Addressing Positivity Violations in Continuous Interventions through Data-Adaptive Strategies and it is similar to the approach described in an earlier proposal 16. These methods allow for the estimation of conditional densities even when the relationship between confounders and interventions is complex or highly non-linear.
@@ 4.1.1 Simulation 1: Correct model specification versus misspecification under positivity violations
We consider a simple scenario with a single time point where both the intervention and the confounder are normal-distributed shown in Appendix Addressing Positivity Violations in Continuous Interventions through Data-Adaptive Strategies. The simulation study involves three variations that share the same distribution for the confounder and intervention but differ in their outcome functions.
@@ 4.1.2 Simulation 2: Correct model specification with positivity violations versus no positivity violations
We consider a different scenario with a single time point where the confounder is binary shown in Appendix Addressing Positivity Violations in Continuous Interventions through Data-Adaptive Strategies. The simulation study involves two variations that share the same distribution for the confounder and but differ in their intervention distribution.
@@ 4.1.3 Simulation 3: Simulation study in more complex setting inspired by real data
In this scenario, we consider a complex data-generating process, detailed in Appendix Addressing Positivity Violations in Continuous Interventions through Data-Adaptive Strategies. This process is inspired by real data from the CHAPAS-3 study, an open-label, parallel-group, randomized trial 17, which is further analyzed in Section 5 and follows the setup described in Schomaker et al. 7.
@@ 4.2 Estimation and Evaluation
4.2.1 Sample Size
For all simulation studies, the sample size is set to $n=1000$.
4.2.2 Causal Parameters
Each estimand is denoted by $m^{\varepsilon}$, where $\varepsilon\in\{\text{standard},\text{feasible},\text{trimming}\}$, as defined in Section 3.1 to 3.3 and its true value is computed using a Monte Carlo simulation with $N=100,000$ samples.
4.2.3 Causal Estimation
For simulation studies 1 and 2, we estimate using the standard parametric g-computation plug-in estimator. For simulation 3, we employed the semi-parametric version. The specific details of the estimation are given in the Algorithm 1.
4.2.4 Conditional Density Estimation
For simulation 1, we performed parametric conditional density estimation via linear regression. In simulation 2, we used non-parametric kernel conditional density estimation, offered in R package np 12. For simulation 3, we implemented super learner based conditional density estimation, as described in Appendix Addressing Positivity Violations in Continuous Interventions through Data-Adaptive Strategies.
4.2.5 Bias Metrics
To evaluate the estimation bias of each estimand, all Monte Carlo simulations are based on $R=1000$ runs.
We define
Absolute $\text{Absolute Bias}(a)=\dfrac{\sum_{r=1}^{R}|\hat{m}^{\varepsilon}_{r}(a)-m^{\varepsilon}(a)|}{R}$
with
$m^{\varepsilon}(a)=\dfrac{\sum_{r=1}^{N}\hat{m}^{\varepsilon}_{r}(a)}{N},$
and $\hat{m}^{\varepsilon}_{r}(a)$ represents the result of each Monte Carlo run.
@@ 4.3.1 Simulation 1
(a) Conditional density distribution of $A|L$ with different coverage level and scatter plot of observed data from one single simulation run.
(b) Non-overlap ratio: the conditional support with coverage level of $95\%$.
@@
Figure 3a illustrates the estimated conditional distribution and coverage levels at $50\%$, $90\%$, $95\%$, and $99\%$, overlaid with a scatter plot of observed data from a single simulation run. Areas outside these regions highlight rare combinations of $L$ and $Y$ in observation, reflecting data sparsity. Using the diagnostic approach from Section 2.3, we identify positivity violations, as shown in Figure 3b.
Figure 3b presents the non-overlap ratio, a diagnostic metric quantifying the proportion of the population for which specific interventions are infeasible at a $95\%$ coverage level. This ratio, ranging from 00 to $1$, reveals a clear pattern: in the central region, nearly all interventions have a non-overlap ratio of zero, indicating good feasibility. In contrast, the ratio rises sharply at the extremes, indicating that extremely small or large interventions are largely infeasible.
@@ 4.3.2 Simulation 2
(b) Non-overlap ratio: the conditional support with coverage level of $95\%$.
@@ 4.3.3 Simulation 3
(a) Non-overlap ratio: the conditional support with coverage level of $95\%$.
@@
Figure 5b presents the true and estimated values for the three estimands. In regions where the non-overlap ratio is zero, the three estimands are indistinguishable. However, the standard estimand increases dramatically near an intervention value of 00. As indicated by the non-overlap ratio, such values are nearly impossible to intervene on. In the data-generating process (see Appendix Addressing Positivity Violations in Continuous Interventions through Data-Adaptive Strategies), the truncated normal distribution excludes values below $0.2032$, causing the true curve of the trimmed estimand to terminate near this intervention value. This indicates that intervention values below $0.2032$ are strictly infeasible, representing a theoretical positivity violations and thus making counterfactual curve of trimming estimand unidentifiable below that threshold. Unlike the standard estimand, the proposed most feasible intervention estimand does not exhibit extreme values in regions with very low intervention, making it more robust and less prone to positivity issues.
@@ 5 Data Analysis
We demonstrate our approach using pharmacoepidemiological data of HIV-positive children from the CHAPAS-3 trial, which enrolled children aged 1 month to 13 years in Zambia and Uganda 17, 18. We focus on those 125 children who received efavirenz as part of antiretroviral therapy, together with both lamivudine and either stavudine, zidovudine, or abacavir.
In line with previous studies 19, 20, 21, we are interested in how counterfactual viral failure probabilities (VLt), defined as a viral load exceeding $100$ copies/mL, vary as a function of different efavirenz concentrations (EFVt, defined as the plasma concentration (in mg/L) measured $12$ hours after dosing). Briefly, different patients who take the same efavirenz dose, may still have different drug concentrations in their blood, for example because of their individual metabolism. Patients exposed to suboptimal concentrations may experience negative outcomes and one may ask at which concentration level antiretroviral activity is too low to guarantee suppression for most patients. Thus, one may be interested in the concentration-response curve $a\mapsto P(\text{VL}^{\text{EFV}=a}=1),\,a\in[0,6]$ mg/L; although it is known that not every child can achieve every possible concentration level, depending on their clinical and metabolic profile and thus positivity violations may be likely 7.
For illustration, we use baseline data and data from the scheduled visit at $t=36$ (weeks). The relevant part of the assumed data-generating process is, in line with previous studies 21, 22, 7, summarized by the directed acyclic graph (DAG) in Figure 6:
@@
Figure 6: Directed acyclic graph (DAG) for CHAPAS-3 trial described in Section 5. NRTI refers to the Nucleoside Reverse Transcriptase Inhibitor drug drug component, sex, age, genotype relates to how fast the drug is cleared as determined by the 516G and 983T polymorphisms in the CYP2B6 gene 19, non-adherence describes any signs of non-adherence as measured through a medication event monitoring system 21 as well as weight. Subscripts describe the the week of measurement.
Our analysis illustrates the ideas based on a complete case analysis of all measured variables at $t=36$ (weeks) represented in the DAG $(n=83)$. The analysis was conducted using three coverage levels for detecting non-overlap: $99\%$, $95\%$, and $90\%$. Our target estimand is the standard causal concentration-response curve, defined in Section 3.1. We compare this estimand to the proposed estimand with most feasible intervention strategy, described in Section 3.2.4, as well as the trimming estimand in Section 3.3. The estimation process closely follows the detailed methodology described in Section 3.4 and algorithm outlined in Algorithm 1. Bootstrap resampling was used to estimate $95\%$ compatibility intervals.
@@ 5.1 Results
(a) Non-overlap ratio at coverage level of $90\%$, $95\%$ and $99\%$. As the coverage level increases, there is greater tolerance for non-overlap detection. Major positivity violations occur in intervention regions with values below $1$ and above $3.5$.
(b) Comparison of different curves including feasible interventions with coverage level $90\%$, $95\%$ and $99\%$, and the standard curve.
(c) Comparison with of different curves. Separated by coverage level of $90\%$, $95\%$ and $99\%$, with $95\%$ bootstrap compatibility intervals.
@@
Figure 7a illustrates the proposed non-overlap ratios as diagnostics for identifying positivity violations in continuous interventions. These ratios, calculated as a function of the intervention $\text{EFV}_{36}$, represent the proportion of the population where the intervention is not feasible, based on 3 coverage levels of non-overlap detection.
@@
The non-overlap ratios remain low across all coverage levels within the central intervention range (approximately $1$ to $3.5$ mg/L), suggesting minimal violations of positivity in this region and supporting reliable causal estimation. However, at the boundaries of the intervention range—below $1$ mg/L or above $3.5$ mg/L—the ratios increase, indicating more frequent violations. Notably, near 00 mg/L, the non-overlap ratio rises sharply, signaling strong violations of the positivity assumption, consistent with the fact that concentrations this low are virtually impossible to observe.
@@
Figure 7b show the estimated curves of the most feasible intervention across three coverage levels compared with the standard causal concentration-response curve. In regions with positivity violations, using the most feasible intervention strategy differs from the CDRC, while closely aligning with standard estimates in well-supported regions. This means whenever a concentration value seems unlikely or implausible, we do not enforce this intervention but rather use the closest value that is possible. This may happen, for example, if there are patients with a very slow metabolism who will never be able to clear the drug fast enough to ever achieve concentrations under $1$ mg/L.
@@
Figure 7c compares all estimated causal curves across three coverage levels. As the coverage level increases, more positivity violations are tolerated for both the feasible intervention and triming estimands. However, the triming estimand exhibits instability in regions with strong positivity violations, suggesting that it may not be the optimal approach in this analysis.
6 Conclusion
Positivity violations are a common issue with continuous interventions, presenting significant challenges when the outcome function cannot be extrapolated reliably from observed to unobserved regions. These violations occur when certain values of the intervention are not adequately represented in the data, leading to a lack of support and, consequently, unreliable estimations.
One of the key consequences of positivity violations is an increase in absolute bias. This issue persists even when the underlying model is correctly specified. In essence, the model’s ability to provide accurate estimates is compromised not only by incorrect specifications but also by the absence of sufficient data in certain regions of the intervention space, which implies addressing positivity violations is not a problem that can be resolved merely by employing robust estimators against misspecification.
Estimating the standard causal dose-response curve (CDRC), an idealized causal representation of the relationship between the intervention and outcome, becomes particularly complex when the range of interventions is extensive. The broad spectrum of interventions can exacerbate the difficulty of making accurate estimations, as it requires the model to reliably estimate across a wide range of values, including those not well-represented in the data.
To address this challenge, we propose using a data-adaptive intervention as a novel strategy to construct a new estimand that tackles positivity issues. The key idea is to adhere as closely as possible to the actual CDRC, deviating only for individuals where the target intervention is clearly infeasible based on the data. In such cases, we substitute the intervention with the most feasible value, ensuring the interpretation remains meaningful: we either apply the intervention of interest or the nearest feasible alternative. By focusing on a more practical and adaptive range of interventions, this approach reduces bias and improves robustness. Consequently, the resulting causal dose-response curve accurately reflects relationships where sufficient data exists, enhancing both precision and interpretability.
In summary, while positivity violations present significant challenges, careful selection of estimands can substantially mitigate variability and enhance the interpretability of the estimates. Future research should explore adapting the concept of the data-adaptive intervention to longitudinal settings, as well as developing and applying different estimators and estimation methods to further refine and validate these approaches. Additionally, there is considerable room for improvement in defining the practical support and refining substitution intervention functions, especially in addressing limitations related to the reliance on conditional density estimation.

### v2

@@ 3.4.2 Conditional Density Estimation
A key step in this process involves estimating the conditional density of the intervention given the confounders. Accurate conditional density estimation is essential for characterizing the regions in which the oracle intervention is supported by the data. Several approaches can be employed, including parametric models, nonparametric kernel conditional density estimators [12], and more flexible methods such as the Highly Adaptive Lasso–based density estimator (Haldensify) [13, 16, 14]. Motivated by Haldensify, in this paper we adopt a hazard-binning–based conditional density estimator, described in Appendix C, which is closely related to the approach of Díaz et al.[10]. This class of estimators can flexibly approximate conditional densities even when the relationship between confounders and interventions is complex or highly nonlinear.
Algorithm 1 Estimation Algorithm for $m^{\text{standard}},m^{\text{feasible}},m^{\text{trimming}}$, and $\tau$
1:Set-Up:
  • Let $n$ denote the number of subjects.
  • Define a grid of $m$ intervention values: $\overline{\mathbf{a}}=(a_{1},\cdots,a_{m})^{\mathrm{T}}$.
  • Choose a support level $\alpha$ to define highest density regions (HDRs).
2:Estimate Highest Density Regions:
  • (a)
    Conditional Density Estimation: Estimate the conditional treatment density $\hat{f}(a\mid\mathbf{l})$ using the observed data, and evaluate $\hat{f}(a_{i}\mid\mathbf{l}_{j})$ for all $i=1,\cdots,m$ and $j=1,\cdots,n$.
  • (b)
    Density Thresholds: For each subject $j$, determine $\hat{f}_{\alpha}^{j}$, the threshold such that the estimated HDR for subject $j$ contains probability mass $\alpha$.
3:Compute the Non-Overlap Ratio:
$\hat{\tau}(a_{i})=\frac{1}{n}\sum_{j=1}^{n}I\big(\hat{f}(a_{i}\mid\mathbf{l}_{j})<\hat{f}_{\alpha}^{j}\big)$
4:Estimate the Outcome Model:
  • Estimate the conditional outcome model $\hat{E}(Y\mid A,\mathbf{L})$, for example using parametric regression.
  • Note: For plug-in estimation and bootstrap inference, we recommend outcome models with established theoretical guarantees (e.g., parametric models). Flexible methods such as HAL can also be considered.
5:Causal Estimation via Plug-in Estimators:
  • (a)
    Standard Estimand:
    $\hat{m}^{\mathrm{standard}}(a_{i})=\frac{1}{n}\sum_{j=1}^{n}\hat{E}(Y\mid A=a_{i},\mathbf{l}_{j})$
  • (b)
    Feasible Estimand:
    • For each $a_{i}$ and subject $j$: if $\hat{f}(a_{i}\mid\mathbf{l}_{j})<\hat{f}_{\alpha}^{j}$, set $\hat{\mathop{}\!\mathrm{d}}(a_{i},\mathbf{l}_{j})$ to the nearest supported value within the estimated HDR; otherwise, set $\hat{\mathop{}\!\mathrm{d}}(a_{i},\mathbf{l}_{j})=a_{i}$.
    • Calculate
      $\hat{m}^{\mathrm{feasible}}(a_{i})=\frac{1}{n}\sum_{j=1}^{n}\hat{E}\big(Y\mid A=\hat{\mathop{}\!\mathrm{d}}(a_{i},\mathbf{l}_{j}),\mathbf{l}_{j}\big)$
  • (c)
    Trimming Estimand:
    • Estimate the trimming function:
      $\hat{t}(a_{i},\mathbf{l}_{j})=\begin{cases}\dfrac{1}{\hat{\tau}(a_{i})},&\text{if }\hat{f}(a_{i}\mid\mathbf{l}_{j})\geq\hat{f}_{\alpha}^{j},\\[6.0pt] 0,&\text{otherwise}.\end{cases}$
    • Calculate
      $\hat{m}^{\mathrm{trimming}}(a_{i})=\frac{1}{n}\sum_{j=1}^{n}\hat{E}(Y\mid A=a_{i},\mathbf{l}_{j})\cdot\hat{t}(a_{i},\mathbf{l}_{j})$
6:Construct Compatibility Intervals (Optional):
  • If using parametric or $\sqrt{n}$-consistent outcome models, and assuming standard regularity conditions, use the nonparametric bootstrap to construct approximate compatibility intervals for $\hat{m}^{\text{standard}}$, $\hat{m}^{\text{feasible}}$, and $\hat{m}^{\text{trimming}}$.
@@ 4.1.1 Simulation 1: Correct model specification versus misspecification under positivity violations
We consider a simple scenario with a single time point where both the intervention and the confounder are normal-distributed shown in Appendix B.1. The simulation study involves three variations that share the same distribution for the confounder and intervention but differ in their outcome functions.
@@ 4.1.2 Simulation 2: Correct model specification with positivity violations versus no positivity violations
We consider a different scenario with a single time point where the confounder is binary shown in Appendix B.2. The simulation study involves two variations that share the same distribution for the confounder and but differ in their intervention distribution.
@@ 4.1.3 Simulation 3: Simulation study in more complex setting inspired by real data
In this scenario, we consider a complex data-generating process, detailed in Appendix B.3. This process is inspired by real data from the CHAPAS-3 study, an open-label, parallel-group, randomized trial [26], which is further analyzed in Section 5 and follows the setup described in Schomaker et al. [31].
@@ 4.2 Estimation and Evaluation
We set the sample size to $n=1000$ for all simulation studies. The true value of each estimand $m^{\varepsilon}$, where $\varepsilon\in\{\text{standard},\text{feasible},\text{trimming}\}$, is computed by intervening on the data-generating process using the specified actions, with a large sample size of N=100, [^0] N=100{,}000. For estimation (see Algorithm 1 for details), we apply the parametric g-computation plug-in estimator in Simulations 1 and 2, and the HAL-based estimator in Simulation 3. The method for conditional density estimation varies by scenario: parametric linear regression (Simulation 1), nonparametric kernel methods (Simulation 2, using the R package np [12]), and Super Learner (Simulation 3; see Appendix C). We assess bias over $R=1000$ Monte Carlo replications, defining absolute bias at level $a$ as:
$\text{Absolute Bias}(a)=\dfrac{1}{R}\sum_{r=1}^{R}|\hat{m}^{\varepsilon}_{r}(a)-m^{\varepsilon}(a)|.$
with $m^{\varepsilon}$ the true target parameter.
@@ 4.3.1 Simulation 1
(a) Conditional density distribution of $A|L$ with different support level and scatter plot of observed data from one single simulation run.
(b) Non-overlap ratio: the conditional support with support level of $95\%$.
@@
Figure 3a illustrates the estimated conditional distribution and support levels at $50\%$, $90\%$, $95\%$, and $99\%$, overlaid with a scatter plot of observed data from a single simulation run. Areas outside these regions highlight rare combinations of $L$ and $Y$ in observation, reflecting data sparsity. Using the diagnostic approach from Section 2.2, we identify positivity violations, as shown in Figure 3b.
Figure 3b presents the non-overlap ratio, a diagnostic metric quantifying the proportion of the population for which specific interventions are infeasible at a $95\%$ support level. This ratio, ranging from 0 to $1$, reveals a clear pattern: in the central region, nearly all interventions have a non-overlap ratio of zero, indicating good feasibility. In contrast, the ratio rises sharply at the extremes, indicating that extremely small or large interventions are largely infeasible.
@@ 4.3.2 Simulation 2
(b) Non-overlap ratio: the conditional support with support level of $95\%$.
@@ 4.3.3 Simulation 3
(a) Non-overlap ratio: the conditional support with support level of $95\%$.
@@
Figure 5b presents the true and estimated values for the three estimands. In regions where the non-overlap ratio is zero, the three estimands are indistinguishable. However, the standard estimand increases dramatically near an intervention value of 0. As indicated by the non-overlap ratio, such values are nearly impossible to intervene on. In the data-generating process (see Appendix B.3), the truncated normal distribution excludes values below $0.2032$, causing the true curve of the trimmed estimand to terminate near this intervention value. This indicates that intervention values below $0.2032$ are strictly infeasible, representing a theoretical positivity violations and thus making counterfactual curve of trimming estimand unidentifiable below that threshold. Unlike the standard estimand, the proposed most feasible intervention estimand does not exhibit extreme values in regions with very low intervention, making it more robust and less prone to positivity issues.
@@ 5 Data Analysis
We demonstrate our approach using pharmacoepidemiological data of HIV-positive children from the CHAPAS-3 trial, which enrolled children aged 1 month to 13 years in Zambia and Uganda [26, 1]. We focus on those 125 children who received efavirenz as part of antiretroviral therapy, together with both lamivudine and either stavudine, zidovudine, or abacavir.
In line with previous studies [4, 5, 30], we are interested in how counterfactual viral failure probabilities (VLt), defined as a viral load exceeding $100$ copies/mL, vary as a function of different efavirenz concentrations (EFVt, defined as the plasma concentration (in mg/L) measured $12$ hours after dosing). Briefly, different patients who take the same efavirenz dose, may still have different drug concentrations in their blood, for example because of their individual metabolism. Patients exposed to suboptimal concentrations may experience negative outcomes and one may ask at which concentration level antiretroviral activity is too low to guarantee suppression for most patients. Thus, one may be interested in the concentration-response curve $a\mapsto P(\text{VL}^{\text{EFV}=a}=1),\,a\in[0,6]$ mg/L; although it is known that not every child can achieve every possible concentration level, depending on their clinical and metabolic profile and thus positivity violations may be likely [30].
For illustration, we use baseline data and data from the scheduled visit at $t=36$ (weeks). The relevant part of the assumed data-generating process is, in line with previous studies [30, 18, 31], summarized by the directed acyclic graph (DAG) in Figure 6:
@@
Figure 6: Directed acyclic graph (DAG) for CHAPAS-3 trial described in Section 5. NRTI refers to the Nucleoside Reverse Transcriptase Inhibitor drug drug component, sex, age, genotype relates to how fast the drug is cleared as determined by the 516G and 983T polymorphisms in the CYP2B6 gene [4], non-adherence describes any signs of non-adherence as measured through a medication event monitoring system [30] as well as weight. Subscripts describe the the week of measurement.
Our analysis illustrates the ideas based on a complete case analysis of all measured variables at $t=36$ (weeks) represented in the DAG $(n=83)$. The analysis was conducted using three support levels for detecting non-overlap: $99\%$, $95\%$, and $90\%$. Our target estimand is the standard causal concentration-response curve, defined in Section 3.1. We compare this estimand to the proposed estimand with most feasible intervention strategy, described in Section 3.3.3, as well as the trimming estimand in Section 3.3.4. The estimation process closely follows the detailed methodology described in Section 3.4 and algorithm outlined in Algorithm 1.
Note: To quantify uncertainty, obtaining reliable compatibility intervals is challenging in this setting, both because the functional parameter defined in Equation (16) is not pathwise differentiable and because of uncertainty around ideal model specification. Since bootstrap inference may be unreliable in such nonregular problems, we report only point estimates in the application and acknowledge this limitation.
@@ 5.1 Results
(a) Non-overlap ratio at support level of $90\%$, $95\%$ and $99\%$. As the support level increases, there is greater tolerance for non-overlap detection. Major positivity violations occur in intervention regions with values below $1$ and above $3.5$.
(b) Comparison of different curves including feasible interventions with support level $90\%$, $95\%$ and $99\%$, and the standard curve.
(c) Comparison with of different curves. Separated by support level of $90\%$, $95\%$ and $99\%$.
@@
Figure 7a illustrates the proposed non-overlap ratios as diagnostics for identifying positivity violations in continuous interventions. These ratios, calculated as a function of the intervention $\text{EFV}_{36}$, represent the proportion of the population where the intervention is not feasible, based on 3 support levels of non-overlap detection.
@@
The non-overlap ratios remain low across all support levels within the central intervention range (approximately $1$ to $3.5$ mg/L), suggesting minimal violations of positivity in this region and supporting reliable causal estimation. However, at the boundaries of the intervention range—below $1$ mg/L or above $3.5$ mg/L—the ratios increase, indicating more frequent violations. Notably, near 0 mg/L, the non-overlap ratio rises sharply, signaling strong violations of the positivity assumption, consistent with the fact that concentrations this low are virtually impossible to observe.
@@
Figure 7b show the estimated curves of the most feasible intervention across three support levels compared with the standard causal concentration-response curve. In regions with positivity violations, using the most feasible intervention strategy differs from the CDRC, while closely aligning with standard estimates in well-supported regions. This means whenever a concentration value seems unlikely or implausible, we do not enforce this intervention but rather use the closest value that is possible. This may happen, for example, if there are patients with a very slow metabolism who will never be able to clear the drug fast enough to ever achieve concentrations under $1$ mg/L.
@@
Figure 7c compares all estimated causal curves across three support levels. As the support level increases, more positivity violations are tolerated for both the feasible intervention and triming estimands. However, the triming estimand exhibits instability in regions with strong positivity violations, suggesting that it may not be the optimal approach in this analysis.
6 Discussion
6.1 Summary
Positivity violations are a persistent challenge in causal inference with continuous interventions, particularly when outcome regressions must be extrapolated from observed to unobserved regions. Such violations arise when certain intervention levels are insufficiently represented in the data, undermining both the identification and estimation of causal effects.
We address this challenge with a data-adaptive framework that employs an $L$-specific density threshold $f_{\alpha}$ to define highest density regions $\mathcal{A}_{\alpha}(\mathbf{l})$. Conceptually, this idea is related to Modified Treatment Policies (MTPs), since both approaches avoid assigning interventions to strata with poor support. The objectives, however, differ: MTPs define modified treatment (often based on natural values) to ensure feasibility, whereas our method preserves the causal dose–response curve (CDRC) as much as possible. As Ring et al. [28] note, even MTPs may not fully resolve positivity violations when the modified treatment lacks adequate support.
Our approach also contrasts with the weighted CDRC proposed by Schomaker et al. [31]. Both approaches address positivity violations via conditional densities, but they differ in how the estimand is adapted. First, we employ a unit-specific, data-adaptive threshold, whereas Schomaker et al. use a fixed cutoff $c$, which may be unsuitable across intervention ranges and covariate strata. Second, our framework serves dual purposes: it functions both as a diagnostic tool (see Section 2.3) and as a procedure for defining a feasible intervention rule. A further distinction lies in interpretability. Our method modifies intervention assignments explicitly (to the “closest feasible” value) and targets the clearly defined quantity $E(Y^{\mathop{}\!\mathrm{d}})$. In contrast, the weighted CDRC redefines the estimand implicitly through a projection within the expectation, thereby changing the target parameter.